Find nature of root of polynomial 2x^5+x^3+5x+1
1
Expert's answer
2020-04-06T16:41:21-0400

Let the given polynomial be "f(x)=2x^5+x^3+5x+1" .

Since it is an odd degree polynomial ,hence it has at least one real roots.

If possible ,"f(x)" have two real roots "a \ and \ b" .

Then "f(a)=f(b)=0" .

Again, since polynomial function are continuous and differential over "\R" "(" In particular over any subset of "\R)"

Therefore "1)f(x)" is continuous at "[a,b]"

"2) f(x)" is differentiable at "(a,b)"

and "f(a)=f(b)=0" .

Then by Rolle's Theorem, there exit at least one point "c\in(a,b)" such that "f'(c)=0" .

But "f'(c)=10x^4+3x^2+5 >0" ,since "x^4\geq0 \ and \ x^2\geq 0" for any real number "x"

Therefore,We get a contradiction.

Hence ,"f(x)" have exactly one real root .

Also the number of imaginary roots"=5-1=4" .

Since ,a polynomial of degree n over a field of complex number have exactly n roots.